OpenStudy (calculus101):

Determine the sum of (3/7)^n+1 from 1 to infinity

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}r^n=\frac{1}{1-r}\]

OpenStudy (calculus101):

\[\sum_{1}^{\infty} (3/7)^\]